Salmon often jump waterfalls to reach their breeding grounds. Starting downstream, 1.9 m away from a waterfall 0.507 m in height, at what minimum speed must a salmon jumping at an angle of 27.9 degrees leave the water to continue upstream? The acceleration due to gravity is 9.81 m/s ^ 2 . Answer in units of m/s .


1
Expert's answer
2021-09-26T18:48:21-0400

The height is


"h=\frac{v^2\sin^2\theta}{2g}.\\\space\\
v=\frac{\sqrt{2gh}}{\sin\theta}=6.74\text{ m/s}."